7=ln(5x)+ln(7x-2x)
Answer by mananth(16949)

7=ln(5x)+ln(7x-2x)
simplify:
7 = ln5x + ln5x
7 = 2ln5x
3.5 = ln5x
Eliminate the ln from the equation raise everything to the e power.
e^3.5 = eln5x
By the identity rule, eln5x is equal to 5x
5x = e^3.5
5x = 33.1
x = 6.62